OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x + y = 8 x − 2y = 2

OpenStudy (anonymous):

x - 2y = 2 x = 2y + 2 4(2y + 2) + y = 8 8y + 8 + y = 8 9y = 0 y = 0 x - 2(0) = 2 x = 2